# Improving Concept Alignment in Vision-Language Concept Bottleneck Models

### v1

@@ Abstract
Concept Bottleneck Models (CBM) map the input image to a high-level human-understandable concept space and then make class predictions based on these concepts. Recent approaches automate the construction of CBM by prompting Large Language Models (LLM) to generate text concepts and then use Vision Language Models (VLM) to obtain concept scores to train a CBM. However, it is desired to build CBMs with concepts defined by human experts instead of LLM generated concepts to make them more trustworthy. In this work, we take a closer inspection on the faithfulness of VLM concept scores for such expert-defined concepts in domains like fine-grain bird species classification and animal classification. Our investigations reveal that frozen VLMs, like CLIP, struggle to correctly associate a concept to the corresponding visual input despite achieving a high classification performance. To address this, we propose a novel Contrastive Semi-Supervised (CSS) learning method which uses a few labeled concept examples to improve concept alignment (activate truthful visual concepts) in CLIP model. Extensive experiments on three benchmark datasets show that our approach substantially increases the concept accuracy and classification accuracy, yet requires only a fraction of the human-annotated concept labels. To further improve the classification performance, we also introduce a new class-level intervention procedure for fine-grain classification problems that identifies the confounding classes and intervenes their concept space to reduce errors. Source codes are available at https://github.com/NMS05/Improving-Concept-Alignment-in-Vision-Language-Concept-Bottleneck-Models.
Keywords:
Concept Bottleneck Model Interpretability Semi-Supervised
1 Introduction
@@
Concept Bottleneck Models (CBM) [14] are a popular choice for building interpretable models with high-level concepts. These models first map the input image to a low-dimensional (hence the term “bottleneck”) expert-defined concept space where each concept is readily understandable by a human. For example, the bird species ‘Vermilion Flycatcher’ can have concepts like ‘red-body’, ‘black-wing’, and ‘red-crown’, etc. A linear classifier then makes class predictions on top of these concepts. There are two fundamental requirements to build a CBM - a well-defined non-ambiguous concept set and the corresponding concept labels for every sample. Recent approaches [25, 46] automate the above steps and scale CBM to larger datasets (like ImageNet) by leveraging Large Language Models (LLM) and Vision Language Models (VLM). They typically involve generating an initial concept set by prompting a LLM with few shot examples and then extracting a sub-set of those concepts that are conducive for discrimination (classification). When an image and this filtered concept set  [^1] The number of LLM generated concepts are usually an order of magnitude larger than the expert-defined concepts. is passed as a input to the contrastively pre-trained VLM, it’s image-text alignment scores act as “concept scores or concept labels” for that sample. These concept labels can be used to train another CBM classifier [25] or can be directly used to predict the classes [46], thus turning a VLM into CBM (referred to as VL-CBM from here on).
@@
While LLMs possess rich world knowledge [27, 11, 34] and can generate a potential (large) concept set across any domain, there are certain instances where an expert-defined concept is preferred.
@@
  • A small and concise set of expert-defined bottleneck concepts enables easier error analysis and intervention, which is crucial for interpretable models.
@@
  • In medical domains, such as Hematology [36, 2], it is desired to use the expert-defined concept set rather than that generated from LLMs, as the former comes from the more trusted pathologists.
  • For faithful “visual” classification, all the concepts must be visually groundable, which is guaranteed with expert-defined concept set [38]. In comparison, for problems like bird species classification, LLMs can generate concepts like “migratory birds” that are not salient for visual classification.
@@
In such cases with limited yet preferable expert concepts, do the VLM concept scores faithfully represent the visual input and are they efficacious to build VL-CBMs with desired performance? To answer this question, we take a closer inspection on the CLIP [28] concept scores for expert-defined concepts on classification datasets like CUB [38], RIVAL [21], and AwA2 [42]. Ideally, it is desired for the CLIP model to precisely associate a visual concept (appearance, color, shape, etc.) to the corresponding visual input. Our investigations expose two problems,
  • Low Concept Accuracy. The CLIP (VLM) model has a low concept accuracy despite the VL-CBM achieving a high classification performance, which makes the resulting model less faithful.
@@
  • Incorrect Concept Association. For challenging classification problems like CUB, the CLIP model is sensitive and biased towards the primary color of the bird and associate this dominant color to all other body parts. It struggles to correctly attribute the fine-grain concepts to the visual input.
@@
Hence, to improve the reliability of VL-CBM, there is a need to improve concept alignment i.e., activate the desired/truthful concepts for a given image, which requires supervision. But obtaining the supervisory concept labels for all training samples is cumbersome. To solve this problem, we propose a novel Contrastive Semi-Supervised (CSS) learning approach that can improve the concept alignment in VL-CBM with fewer labeled concept examples. Our approach encourages consistent concept scores for intra-class samples whilst discriminating (contrasting) them from the other classes. It then uses a few labelled concept examples (semi-supervision) per class to align them with the ground truth. Extensive experiments on three benchmark datasets CUB [38], RIVAL [21], and AwA2 [42] show that our CSS method substantially increases the concept accuracy (+39.1% for CUB, +18.63% for RIVAL, +31.11% for AwA2) and enhances the overall classification accuracy (+5.61% for CUB, +2.84% for RIVAL, +3.06% for AwA2) with only a handful of human-annotated concept labels per class. We also validate our CSS approach in a medical domain using the blood cell attribute dataset WBCAtt [36].
@@
Intervention is the fundamental advantage of a CBM over other interpretable models - if a classifier makes an incorrect prediction, one can inspect the concept scores and rectify the problem, so that the model makes the desired prediction. Intervention in CBM is typically performed at an instance level [14, 25], where a handful of error images are arbitrarily chosen for debugging. But, choosing the error images are non-trivial for fine-grain multi-class classification problems which have fewer samples per class and even visually similar classes. In this work, we propose a class-level intervention procedure for fine-grain classification problems, where we first compute an error matrix to identify the “confounding classes” (visually similar but semantically different) and then intervene the error images of these classes such that it reduces the total error for the confounding classes whilst increasing the overall classification accuracy. The primary contributions of this paper are summarized below,
  • Validating the faithfulness of VL-CBM. We investigate the concept scores for expert concepts and show that CLIP models have poor concept alignment and they struggle to correctly associate the fine-grain concepts to visual input.
@@
  • Improving concept alignment. To address the above issue and to improve the reliability of VL-CBM, we propose a novel Contrastive Semi-Supervised (CSS) approach which uses a fraction of the human-annotated concept labels to increase the concept accuracy and classification accuracy.
@@
  • Class-level Intervention. We propose an intervention procedure for fine-grain classification problems which reduces the total error for confounding classes and further increases the classification performance.
2 Related Work
Post-hoc interpretable models. Post-hoc methods [30, 29, 17, 4, 8] convert a pre-trained black-box classifier into an interpretable model using saliency maps. They highlight the regions of the input space [31] that influence a classifiers’ prediction and it helps one understand the salient features of a particular class. Though saliency maps can reveal the semantic features (or spurious correlations) that a classifier focuses on, their interpretation can be subjective and may require domain expertise.
@@ 2 Related Work
Prototype-based and Concept-based interpretable models. In stark contrast to post-hoc methods, these models can be trained and are interpretable by design. Prototype-based methods [6, 24, 32, 37, 7] learn interpretable visual feature prototypes for every class from the training samples. During inference, the distance (similarity score) between the test sample feature and the learned prototypes determine the classifiers decision. BotCL [39] is a prototype-based method that is closely related to our work. It applies contrastive learning to the prototype space of a vision-only model, while we focus on contrastive + semi-supervised learning on the concept space of a vision-language (multimodal) model. Concept-based methods [12, 14, 48] map the input to a high-level concept space defined by human experts, where these concepts are readily interpretable. The classifiers decision is based on the predicted concept scores. The main advantage of concept-based methods is that they allow test-time intervention [14, 12, 19, 1, 5] and they are easier to debug compared to other methods.
@@
VLM based interpretable models (VL-CBM). The problem with concept-based models is that they require concept labels for every training sample [14] or a lot of training images representing each of those concepts [12, 48]. To solve this problem, recent approaches [25, 46, 49, 45, 41, 22, 23] leverage contrastively pre-trained Vision-Language Models, like CLIP [28], to automatically generate concept labels from their image-text alignment scores. These models have interpretable natural language concepts [18, 44] that are typically generated by a LLM and this automated approach scales them to larger datasets.
Enhancing faithfulness of VL-CBMs. Several recent works have focused on addressing the shortcomings of VLM based concept bottlenecks to improve their reliability. They constrain the class predictions to focus on a sub-set of salient concepts [40, 33, 26], make concept scores robust to perturbations [15], and obtain reliable concept score distributions by modeling uncertainties [13] and by composing energy functions [43]. In our work, we enhance the faithfulness of VL-CBM by using human (expert) preferred concepts and improve the concept scores of VLM in a semi-supervised setup using a few concept labels.
3 Methodology
We formally define VL-CBM in Sec. 3.1 and investigate the faithfulness of CLIP concept scores in Sec. 3.2. We then propose our Contrastive Semi-Supervised (CSS) method to improve concept alignment in Sec. 3.3.
3.1 Vision-Language Concept Bottleneck Models
@@
Consider a dataset $D=\{(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{N},y_{N})\}$, where $x_{i}\in\mathbb{R}^{H\times W\times 3}$ denotes the images, $y_{i}$ denotes the labels for $k$ classes, and $N$ is the total number of training samples. While a black-box classifier $f:x\rightarrow\mathbb{R}^{k}$ directly maps an image $x$ to its label $y$, a CBM consists of a backbone $g:x\rightarrow\mathbb{R}^{c}$ which first maps the image $x$ to a concept space consisting of $c$ pre-defined concepts and a linear classifier $h:\mathbb{R}^{c}\rightarrow\mathbb{R}^{k}$ that maps the predicted concepts to its corresponding label $y$. These concepts represent the salient visual features that are crucial for solving the classification task.
@@
However, training the backbone $g$ requires $N\times c$ concept labels which is tedious to obtain. VL-CBM overcome this issue by leveraging the image-text alignment scores of contrastively pre-trained Vision-Language Models (like CLIP) as concept labels. Concretely, let $T=\{t_{1},t_{2},...,t_{c}\}$ be the set of LLM generated or expert-defined text concepts. Now the $c$ concept labels for an image $x_{i}$, which is denoted as $C_{i}$, is obtained by $C_{i}=E_{I}(x_{i})\cdot E_{T}(T)^{\intercal}$, where $\cdot$ denotes the dot product, $E_{I}$ and $E_{T}$ denotes the image and text encoder of the CLIP model that maps the images and text concepts into a shared feature space. Now, with the availability of image $x_{i}$, concept label $C_{i}$, and class label $y_{i}$ for $N$ samples, the backbone $g$ and the classifier $h$ can be independently or jointly trained.
Table 1: Evaluation of CLIP concept scores and VL-CBM classification performance.
@@
   CUB    RIVAL    AwA2 Classification Accuracy (%)    75.84    95.63    90.14 Concept Accuracy (%)    24.43    58.85    49.02 No.of Concepts    312    18    85 Difficulty of Concepts    Hard    Easy    Medium
3.2 Faithfulness of Vision-Language Concept Bottleneck Models
@@
We validate the faithfulness of VL-CBM for expert concepts across three image classification datasets namely Caltech-UCSD Birds (CUB) [38], Rich Visual Attributes with Localization (RIVAL) [21], and Animal with Attributes (AwA2) [42]. These datasets provide expert concepts and the corresponding ground-truth concept labels required for our evaluation. We input the images and the expert (text) concepts to a frozen CLIP model to obtain the concept scores as described in Sec. 3.1. Our objective here is to evaluate the quality and faithfulness of the VLM concept scores by comparing them against the ground-truth concept labels and evaluate their classification performance with a linear classifier. The concept evaluation results are presented in Table 1. Our investigations expose two problems with CLIP concept scores.
@@
Low Concept Accuracy. From the table, the CLIP model has a low concept accuracy despite the VL-CBM achieving a high classification accuracy. This implies that a higher classification performance does not guarantee a faithful concept understanding, which makes the VL-CBM less reliable. Moreover, it becomes difficult to intervene such a model during test-time, in order to rectify prediction errors.
@@
Incorrect Concept Association. A closer inspection of these concepts in Fig. 2 (CLIP Concept Scores) reveal that, for fine-grained concepts like in CUB, the CLIP model is sensitive and biased towards the primary color of the bird and associate this dominant color to all other body parts. The untrained model struggles to correctly associate the color/texture concepts to the corresponding bird part, which is crucial for bird species classification.
The lower concept accuracy and the errors in concept association is a result of the poor relational and compositional understanding of CLIP models i.e., they make errors in associating objects to their attributes [47], especially in a multi-object and multi-attribute setting [16]. Hence, to improve the reliability and faithfulness of VL-CBM, there is a need to improve concept alignment i.e. activate the truthful concepts for the given image.
@@
Figure 1: (a) Architecture of VL-CBM with concept projection (bootstrap). (b) Overview of the Contrastive Semi-Supervised (CSS) learning method.
3.3 Improving concept alignment in VL-CBM
We propose a novel Contrastive Semi-Supervised (CSS) learning method that can improve the concept alignment in VL-CBM with a few labelled concept examples per class. The overall framework is illustrated in Fig. 1.
@@ 3.3 Improving concept alignment in VL-CBM
From Table 1, we can see that the frozen CLIP model achieves a low yet non-trivial concept accuracy. We bootstrap this by adding a learnable linear concept projection layer (see Fig. 1a) that enhances the CLIP concept scores by directly predicting the $c$ concepts from the average pooled visual patch tokens as,
@@
Now the model must be optimized to improve concept alignment and achieve high classification accuracy. To improve concept alignment, the predicted concepts can be compared with the ground truth concept labels. However, the availability of fewer concept labels (semi-supervision) makes it less effective in improving concept alignment across other unlabelled samples. To mitigate this, we augment a contrastive learning objective to the concept space with the following intuition. The contrastive objective encourages “consistent” concept scores for intra-class samples and the concept labels for a few of these samples (semi-supervision) guide or align the predicted concepts with the ground truth. Furthermore, contrastive learning discriminates the concept scores of inter-class samples thus aiding in classification. Such a learning objective of enforcing consistency between sample pairs guided by a few labeled examples [3, 20] has not been studied in VL-CBMs before.
To achieve this (see Fig. 1b), we perform pairwise sampling such that each training sample consists of a pair of images $(x_{i},x_{j})$ from the same class and every other sample in the mini-batch belongs to different classes. There are $n$ samples ($2n$ images) in each mini-batch. For these image pairs, we compute the concept scores $(C_{i},C_{j})$ and class predictions $K=[K_{i},K_{j}]~{}$ [^2] The square brackets denote the tensors are stacked across batch dimension and no longer need to be in pairs.. Let $g=[g_{i},g_{j}]$ and $y=[y_{i},y_{j}]$ be the corresponding ground-truth concept labels and class labels. We jointly train the concept projection layer and the linear classifier to minimize the following loss objective.
[…]
@@
where $sim$ is the cosine similarity metric, $\tau$ is the temperature parameter, $L_{ce}$ and $L1$ are the standard cross entropy loss and mean absolute error, $s$ is the softmax normalization and $\gamma$ is a constant that increases the magnitude of the $L1$ loss function. $l$ denotes the index of the concept score, class logit, and label tensor within the mini-batch, after stacking along the batch dimension.
@@
4 Experiments
4.1 Experimental Setup
Dataset. We validate the effectiveness of our Contrastive Semi-Supervised (CSS) method in improving concept alignment on the CUB [38], RIVAL [21], AwA2 [42] datasets, and in a medical setting on the White Blood Cell Attribute (WBCAtt) dataset [36]. All the above datasets provide an expert-defined concept set and the corresponding ground-truth concept labels, which is crucial for our evaluation.
Model and Training. We use the PLIP ViT-B-16 [9] model available in Hugging Face for experiments on WBCAtt dataset and use the OpenCLIP ViT-B-16 model [10] for experiments on all other datasets. We use Adam optimizer to train all our models and evaluate them with classification accuracy and/or concept accuracy metrics.
@@ 4.1 Experimental Setup
Model CUB RIVAL AwA2 WBCAtt  Class  Concept  Class  Concept  Class  Concept Attribute ViT-B/16 (Black Box)  79.21 NA  99.47 NA   94 NA  66.11 CLIP VL-CBM [25, 46]  75.84 24.43  95.63 58.85  90.14 49.02 42.56 CSS VL-CBM (Ours)  81.45 63.53  98.47 77.48   93.2  81.13  64.19
4.2 Improving Concept Alignment
@@
Bird and Animal classification. We validate our CSS method with concept accuracy and classification accuracy metrics for CUB, RIVAL, and AwA2 datasets. We first train a black-box image classifier with frozen CLIP vision encoder (ViT-B/16) as the backbone to establish the baseline classification performance. The black-box classifier directly predicts the classes from the image features without any concept bottlenecks. We then train a CLIP VL-CBM classifier with expert-defined text concepts as described in Sec. 3.1. From the table, we can see that the classification performance of CLIP VL-CBM is on-par with the black-box classifier, however, it has a lower concept accuracy.

### v2

@@ Abstract
Concept Bottleneck Models (CBM) map images to human-interpretable concepts before making class predictions. Recent approaches automate CBM construction by prompting Large Language Models (LLMs) to generate text concepts and employing Vision Language Models (VLMs) to score these concepts for CBM training. However, it is desired to build CBMs with concepts defined by human experts rather than LLM-generated ones to make them more trustworthy. In this work, we closely examine the faithfulness of VLM concept scores for such expert-defined concepts in domains like fine-grained bird species and animal classification. Our investigations reveal that VLMs like CLIP often struggle to correctly associate a concept with the corresponding visual input, despite achieving a high classification performance. This misalignment renders the resulting models difficult to interpret and less reliable. To address this issue, we propose a novel Contrastive Semi-Supervised (CSS) learning method that leverages a few labeled concept samples to activate truthful visual concepts and improve concept alignment in the CLIP model. Extensive experiments on three benchmark datasets demonstrate that our method significantly enhances both concept ($+29.95\%$) and classification ($+3.84\%$) accuracies yet requires only a fraction of human-annotated concept labels. To further improve the classification performance, we introduce a class-level intervention procedure for fine-grained classification problems that identifies the confounding classes and intervenes in their concept space to reduce errors. Source codes are available at https://github.com/NMS05/Improving-Concept-Alignment-in-Vision-Language-Concept-Bottleneck-Models.
Introduction
@@
Concept Bottleneck Models (CBM) (Koh et al. 2020) are a popular choice to build interpretable models using human-understandable concepts. These models first map input images to a low-dimensional, expert-defined concept space. For instance, the bird species ‘Vermilion Flycatcher’ can be described using concepts like ‘red-body’ and ‘black-wing’. A linear classifier then makes predictions based on these concepts. Building a CBM requires a well-defined, non-ambiguous set of concepts and the corresponding concept labels for each sample, which is labor-intensive to obtain. Recent approaches (Oikarinen et al. 2023; Yang et al. 2023) circumvent this issue by automating CBM construction using Large Language Models (LLMs) and Vision Language Models (VLMs). These methods generate an initial concept set by prompting an LLM with few-shot examples, then extracting a subset of those concepts that are conducive to classification. When an image and this filtered concept set are input to a contrastively pre-trained VLM, its image-text alignment scores serve as “concept scores” or “concept labels” for that sample. These scores can be used to train another CBM classifier or directly predict the classes, effectively turning a VLM into a CBM (referred to as VL-CBM hereon).
@@
While LLMs possess rich world knowledge (Petroni et al. 2019; Jiang et al. 2020) and can generate extensive concept sets across various domains, expert-defined concepts are sometimes preferred for several reasons.
@@
  • Concise Concept Sets: A small, well-defined set of expert concepts allows for easier error analysis and intervention, which is crucial for interpretable models.
@@
  • Trusted Expertise: In medical domains like Hematology (Tsutsui, Pang, and Wen 2023), expert concepts are preferred over LLM-generated ones, as the former comes from the more trusted pathologists.
  • Visual Grounding: For faithful visual classification, all concepts must be visually grounded in the image, which is ensured with expert concepts.
@@
In such cases with limited yet preferable expert concepts, it is essential to ask: Do VLM concept scores faithfully represent the visual input, and are they effective for building VL-CBMs with the desired classification performance? To answer this, we closely examine CLIP (Radford et al. 2021) concept scores for expert concepts on classification datasets such as CUB (Wah et al. 2011), RIVAL (Moayeri et al. 2022), and AwA2 (Xian et al. 2018). Ideally, the CLIP model should accurately associate visual concepts (appearance, color, shape, etc.) with the corresponding visual input. However, our investigations reveal two key issues:
  • Poor Concept Alignment: The CLIP model exhibits relatively lower concept accuracy despite the higher classification performance of VL-CBMs. This suggests that CLIP’s concept scores do not faithfully represent the visual input, thus making the resulting model less reliable.
@@
  • Challenges with Fine-Grained Concepts: For challenging classification tasks like CUB, the CLIP model tends to be biased towards the primary color of the bird, incorrectly attributing this dominant color to other body parts. It struggles to correctly attribute the fine-grain concepts to the visual input.
@@
Hence, to improve the faithfulness and reliability of VL-CBMs, there is a need to improve concept alignment i.e., activate the desired/truthful concepts for a given image, which requires concept supervision. However, obtaining supervisory concept labels for all training samples is cumbersome. To address this, we propose a novel Contrastive Semi-Supervised (CSS) learning approach that improves concept alignment in VL-CBM with fewer labeled concept examples. Our approach encourages consistent concept scores for intra-class samples while contrasting them with other classes. It then aligns these scores with ground truth using a few labeled examples per class (semi-supervision). Extensive experiments on three benchmark datasets CUB, RIVAL, and AwA2 show that our CSS method substantially increases the concept accuracy (+39.1% for CUB, +18.63% for RIVAL, +32.11% for AwA2) and enhances the overall classification accuracy (+5.61% for CUB, +2.84% for RIVAL, +3.06% for AwA2) with minimal human-annotated concept labels. We also validate the effectiveness our CSS approach in a medical domain using the blood cell attribute dataset WBCAtt (Tsutsui, Pang, and Wen 2023).
@@
Intervention is the key advantage of a CBM over other interpretable models, where incorrect predictions of the classifier can be rectified by examining and modifying the concepts. In this work, we propose a class-level intervention procedure for fine-grain classification problems, where we first compute an error matrix to identify the “confounding classes” (visually similar but semantically different) for which the model makes the most error. We then intervene the error images of these classes to improve overall classification accuracy. The primary contributions of this paper are summarized below,
  • Evaluating faithfulness of VL-CBMs. We quantify and evaluate the faithfulness of VLM concept scores, revealing that CLIP models exhibit poor concept alignment and struggle with fine-grain concept association.
@@
  • Improving Concept Alignment: We introduce a novel Contrastive Semi-Supervised (CSS) approach that leverages a fraction of human-annotated and expert-defined concept labels, which are inherently more reliable, to improve both concept and classification accuracies.
@@
  • Class-level Intervention: We propose an intervention procedure for fine-grain classification tasks that reduces errors in confounding classes, further enhancing classification performance.
Related Work
@@
Early works on Explainable AI focused on unimodal (image) classifiers and they can be broadly classified into two categories. Prototype-based methods (Chen et al. 2019; Nauta, Van Bree, and Seifert 2021; Donnelly, Barnett, and Chen 2022; Wang et al. 2023) learn interpretable visual feature prototypes for every class from the training samples. During inference, the distance (similarity score) between the test sample feature and the learned prototypes determines the classifier’s decision. Concept-based methods (Kim et al. 2018; Koh et al. 2020; Zarlenga et al. 2022) map the input to a high-level concept space defined by human experts, where these concepts are readily interpretable. The classifier decision is then based on the predicted concept scores. The main advantage of concept-based methods is that they allow test-time intervention (Mitchell et al. 2021; Abid, Yuksekgonul, and Zou 2021; Bontempelli et al. 2023) and they are easier to debug compared to other methods.
@@
Vision-Language interpretable models (VL-CBMs). A limitation of concept-based models is that they require concept labels for every training sample (Koh et al. 2020) or a lot of training images representing each of those concepts (Kim et al. 2018). To mitigate this, recent approaches (Oikarinen et al. 2023; Yang et al. 2023; Yan et al. 2023b) leverage contrastively pre-trained Vision-Language Models, like CLIP, to automatically generate concept labels from their image-text alignment scores. These models have natural language concepts (Menon and Vondrick 2023; Yan et al. 2023a) which make them easily interpretable.
Enhancing faithfulness. Though pioneering works on VL-CBMs have proposed methods to extract salient concepts for higher classification accuracy, recent studies have identified several drawbacks, such as concept prompts being biased towards object names (Zang et al. 2024), concepts being non-factual or irrelevant (Fang et al. 2024), and insufficient concept coverage (Shang et al. 2024). These recent works improve VL-CBM reliability by discovering concepts that better represent visual attributes (Esfandiarpoor, Menghini, and Bach 2024; Chiquier, Mall, and Vondrick 2024) and have stronger CLIP alignment scores. In comparison, we use expert concepts, which are inherently more trustworthy but limited in size. Importantly, our CSS method can be applied to these works provided a few labels can be annotated for those concepts. Some works also enhance VL-CBMs’ reliability by making the concepts resilient to perturbations (Lai et al. 2023) and using sparse classifiers (Panousis, Ienco, and Marcos 2023; Semenov et al. 2024).
Methodology
Vision-Language Concept Bottleneck Models
@@
We first formally define VL-CBMs. Consider a dataset $D=\{(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{N},y_{N})\}$, where $x_{i}\in\mathbb{R}^{H\times W\times 3}$ denotes the images, $y_{i}$ denotes the labels for $k$ classes, and $N$ is the total number of training samples. While a black-box classifier $f:x\rightarrow\mathbb{R}^{k}$ directly maps an image $x$ to its label $y$, a CBM consists of a backbone $g:x\rightarrow\mathbb{R}^{c}$ which first maps the image $x$ to a concept space consisting of $c$ pre-defined concepts and a linear classifier $h:\mathbb{R}^{c}\rightarrow\mathbb{R}^{k}$ that maps the predicted concepts to its corresponding label $y$. These concepts represent the salient visual features crucial for solving the classification task and are human interpretable.
@@
However, training the backbone $g$ requires $N\times c$ concept labels which is tedious to obtain. VL-CBMs overcome this issue by leveraging the image-text alignment scores of contrastively pre-trained VLMs (like CLIP) as concept scores. Concretely, let $T=\{t_{1},t_{2},...,t_{c}\}$ be the set of LLM generated or expert-defined text concepts. Now the $c$ concept scores for an image $x_{i}$, which is denoted as $C_{i}$, is obtained by $C_{i}=E_{I}(x_{i})\cdot E_{T}(T)^{\intercal}$, where $\cdot$ denotes the dot product, $E_{I}$ and $E_{T}$ denotes the image and text encoder of the CLIP model that maps the images and text concepts into a shared feature space. Now, with the availability of image $x_{i}$, concept score $C_{i}$, and class label $y_{i}$ for $N$ samples, the linear classifier $h$ can be trained to make class predictions.
@@
   CUB    RIVAL    AwA2 Classification Acc (%)   75.84   95.63   90.14 Concept Acc (%)    24.43    58.85    49.02 No.of Concepts    312    18    85 Concept Difficulty    Hard    Easy    Medium
Table 1: Evaluating faithfulness of CLIP concept scores and VL-CBM classification performance.
Evaluating Faithfulness of VL-CBMs
@@
The trustworthiness and reliability of a VL-CBM depends on the faithfulness of its concept scores, which serve two purposes: providing rich representations for a linear classifier and enabling human interpretation of the classifier’s predictions. Prior works on VL-CBMs have focused primarily on classification accuracy, which measures how well samples are separable in the concept space, but doesn’t measure the faithfulness of concept scores in representing visual truth. Therefore, we evaluate both the faithfulness of VLM concept scores by comparing them to ground-truth concept labels (Koh et al. 2020) and their classification performance with a linear classifier. We use three image classification datasets - Caltech-UCSD Birds (CUB) (Wah et al. 2011), Rich Visual Attributes with Localization (RIVAL) (Moayeri et al. 2022), and Animal with Attributes (AwA2) (Xian et al. 2018) because they provide expert-defined concepts and the corresponding ground-truth concept labels needed for our evaluation. The results are presented in Table 1 and a visualization of these concepts is shown in Figure 2. Our investigations reveal two key issues with CLIP concept scores.
@@
Poor Concept Alignment: From the table, the CLIP model has relatively lower concept accuracy across all three datasets, despite the VL-CBM achieving high classification accuracy. This indicates that CLIP’s concept scores do not faithfully represent the visual input and a higher classification performance does not guarantee a faithful concept understanding. Consequently, the model is less reliable, and it becomes difficult to intervene such a model during test-time.
@@
Challenges with Fine-Grained Concepts: A closer inspection of these concepts in Figure 2 (CLIP Concept Scores) reveals that, for challenging tasks like CUB, the CLIP model tends to be biased towards the primary color of the bird and incorrectly attributes this dominant color to other body parts. It struggles to correctly associate the compositional and fine-grain concepts with the visual input.
Why does CLIP exhibit poor concept alignment? Although the CLIP model excels at associating images with text containing object names and attributes, achieving strong zero-shot performance on open-ended classification tasks, we find that the CLIP model struggles with fine-grained visual understanding when only concept descriptions are presented, making errors in associating objects with their attributes. The CLIP model shows unsatisfactory concept accuracy when dealing with such fine-grained concepts defined by experts. Similar findings are also discussed in some previous works (Zang et al. 2024; Pham et al. 2024; Yuksekgonul et al. 2022; Lewis et al. 2022; Yun et al. 2022). As poor concept alignment can reduce the reliability and interpretability of VL-CBMs, it is critical to improve the faithfulness of VL-CBMs by activating the truthful concepts for the given image.
@@
Figure 1: (a) Modified CLIP (VLM) with concept projection layer which enhances the raw concept scores. (b) Architecture of VL-CBM and an overview of the Contrastive Semi-Supervised (CSS) learning method.
Enhancing Faithfulness of VL-CBMs
@@
To improve concept alignment in VL-CBM we propose a novel Contrastive Semi-Supervised (CSS) learning method that leverages a few labelled concept examples per class. Figure 1 illustrates the overall framework. As shown in Table 1, the frozen CLIP model achieves a low yet non-trivial concept accuracy. We enhance this by adding a learnable linear concept projection layer (see Figure 1a) which improves the CLIP concept scores by directly predicting $c$ concepts from the average pooled visual patch tokens as,
@@
Now, the model must be optimized to enhance both concept and classification accuracies. To improve concept accuracy, the concept projection layer can be trained with a few concept labels (semi-supervision), which refines the weak concept scores of CLIP and aligns the predicted concepts with the ground truth. However, one needs to effectively leverage this sparse (concept) supervisory signals and extend them to other unlabelled examples. Contemporary unimodal works (Hu et al. 2024) compare the features of labelled and unlabelled examples in the cosine space, and then assign pseudo concept labels for unlabelled examples using KNN algorithm. But, this necessitates additional computation steps (KNN) for each mini-batch. We instead propose to bootstrap the sparsely supervised concept scores to other unlabelled samples using contrastive learning. This method promotes consistent concept scores for samples sharing ‘similar’ concepts (positive pairs) and discriminates concept scores for samples with ‘disjoint’ concepts (negative pairs). For all datasets in our paper, the similar concept pairs come from intra-class samples and the disjoint concept pairs come from inter-class samples. Furthermore, discriminating the scores of inter-class samples make classes more separable in the concept space, thus aiding in classification.
To achieve this (see Figure 1b), we perform pairwise sampling such that each training sample consists of a pair of images $(x_{i},x_{j})$ from the same class, while every other sample in the mini-batch belongs to different classes. There are $n$ samples ($2n$ images) in each mini-batch, and only a subset of these images will be concept supervised. For these image pairs, we compute the concept scores $(C_{i},C_{j})$ and class predictions $K=[K_{i},K_{j}]~{}$ [^1] The square brackets denote the tensors are stacked across batch dimension and no longer need to be in pairs.. Let $g=[g_{i},g_{j}]$ and $y=[y_{i},y_{j}]$ be the corresponding ground-truth concept labels and class labels. We jointly train the concept projection layer and the linear classifier to minimize the following loss objective.
@@
where $sim$ is the cosine similarity metric, $\tau$ is the temperature parameter, $L_{ce}$ and $L1$ is the standard cross entropy loss and mean absolute error, $s$ is the softmax normalization and $\gamma$ is a constant that increases the magnitude of the $L1$ loss function. $G_{l}$ represents the ground truth concept label for the $l$-th entry. $l$ denotes the index of the concept score, class logit, and label tensor within the mini-batch, after stacking along the batch dimension.
@@
Model CUB RIVAL AwA2 WBCAtt  Class  Concept  Class  Concept  Class  Concept Attribute ViT-B/16 (Black Box)  79.21 NA  99.47 NA   94 NA  66.11 CLIP VL-CBM  75.84 24.43  95.63 58.85  90.14 49.02 42.56 CSS VL-CBM (Ours)  81.45 63.53  98.47 77.48   93.2  81.13  64.19
@@
Experiments
Experimental Setting
Dataset. We validate the effectiveness of our Contrastive Semi-Supervised (CSS) method in improving concept alignment on the CUB (Wah et al. 2011), RIVAL (Moayeri et al. 2022), AwA2 (Xian et al. 2018) datasets, and in a medical setting on the White Blood Cell Attribute (WBCAtt) dataset (Tsutsui, Pang, and Wen 2023). All the above datasets provide an expert-defined concept set and the corresponding ground-truth concept labels, which is crucial for our evaluation.
Model. We use the PLIP ViT-B-16 (Huang et al. 2023) model available in Hugging Face for experiments on WBCAtt dataset and use the OpenCLIP ViT-B-16 model (Ilharco et al. 2021) for experiments on all other datasets, unless otherwise specified.
Training and Evaluation. Following the literature (Koh et al. 2020; Fong and Vedaldi 2017), we use concept accuracy to quantify concept alignment. We use Adam optimizer to train all our models and evaluate them with classification accuracy and concept accuracy metrics. We also conduct a distributional analysis of the VL-CBM concept scores to show the concept alignment explicitly.
Improving Concept Alignment
@@
Bird and Animal classification. We validate our CSS method with concept accuracy and classification accuracy metrics for the CUB, RIVAL, and AwA2 datasets. We first train a black-box image classifier with frozen CLIP vision encoder (ViT-B/16) as the backbone to establish baseline classification performance. The black-box classifier directly predicts the classes from image features without any concept bottlenecks. We then train a CLIP VL-CBM (Oikarinen et al. 2023; Wang et al. 2023) classifier with expert-defined text concepts. From the table, we can see that the classification performance of CLIP VL-CBM is on par with the black-box classifier, however, it has a lower concept accuracy.